OpenStudy (anonymous):

6. –3, –12, –48, ... (5 points)–192; –768; –3,072; geometric; Start with –3 and multiply by 4 repeatedly. –198; –788; –3,072; arithmetic; Start with 4 and multiply by –3 repeatedly. –188; –768; –3,058; arithmetic; Start with 4 and multiply by –3 repeatedly. –194; –788; –3,072; geometric; Start with 3 and multiply by 4 repeatedly

OpenStudy (anonymous):

Tell whether each sequence is arithmetic, geometric, or neither. Find the next three terms of each sequence. Write a rule to describe the sequence

OpenStudy (anonymous):

THis is a geometric series because there is a common ratio rather than a common difference. Then, multiply the next terms by 4 each. -48*4 = -192*4 = -768*4 = -3072 -192 -768 -3072 Multiply by 4 repeatedly.